OpenStudy (anonymous):

Rewrite the slope-intercept form equation into standard form. y+5/2

OpenStudy (anonymous):

i mean y=5/2

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (shoba):

y = mx + c is the standard form... 0 is the slope y = 0x + 5/2

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

A. 2x = 5 B. x + 2y = 5 C. 2y = 5 D. 5y = 2

OpenStudy (shoba):

D is the answer.... 0x - y =5/2 (multiple both sides by 2) 2y = 5

OpenStudy (anonymous):

thank you can you help me with another

OpenStudy (shoba):

y = 5/2 In true slope intercept form (y=mx +b): y = 0x + 5/2 where 0 is the slope and 5/2 is the y-intercept In standard form(ax+by=c): 0x + 1y = 5/2 multiply both sides by 2 0x + 2y = 5 where a =0, b = 2 and c = 5

OpenStudy (shoba):

yes sure